1.1For the X-value 1 the Y-value is

1.2 for the X-value the Y-value is 17

1.3 with how many does the y-value increase for x-value

1.4 at 0 on the x-axis the y-value is 1.explain

1.5 why does the graph form a straight (not a curved line)


1
Expert's answer
2021-07-19T18:37:50-0400

Equation : "x-y+1 = 0"


Graph:





(1.1) For x = 1

Put x=1 in above equation , we get

"x-y+1=0\\1-y+1=0\\y=2"

Hence, for x =1 , y = 2



(1.2) For y = 17

Put y = 17 in above equation , we get

"x-y+1=0\\x-17+1=0\\x=16"

Hence , for x = 16 ; y = 17



(1.3) From the above graph , we clearly say that for all values of "x\in R" , the value of y increases.


(1.4) From the equation,

"x-y+1=0"

For x =0

"0-y+1=0\\y=1"

So, the graph will cut the y- axis at +1 for x =0

(1.5) Since, the equation is linear equation

and the y value changes by the same amount each time as we change the x value.